classify 3y=4x-1 as inconsistent, dependent, or independent
x=4/3y

Answer by stanbon(48558)

classify 3y=4x-1 as inconsistent, dependent, or independent
x=4/3y
-------------
Rearrange:
y = (4/3)x-(1/3)
y = (3/4)x
---------------
The slopes are different so the system is independent.
===========================
Cheers,
Stan H.